OpenStudy (anonymous):

7x - 2 = 2y 3x = 2y - 1 Solve the system of equations by substitution. Answers: (3/4, 13/8) (1/4, -1/8) no solution coincident I got conicident? yes or no?

OpenStudy (solomonzelman):

do you HAVE to use substitution? Elimination would honestly be much better.

OpenStudy (anonymous):

you think?

OpenStudy (solomonzelman):

yes, of course

OpenStudy (solomonzelman):

\(\large\color{slate}{ 7x - 2 = 2y }\) \(\large\color{slate}{ 3x = 2y - 1 }\) rearranging the 2nd equation \(\large\color{royalblue}{ 7x - 2 = 2y }\) \(\large\color{royalblue}{ 3x +1= 2y }\) multiplying the 2nd equation by -1 \(\large\color{royalblue}{ 7x - 2 = 2y }\) \(\large\color{royalblue}{ -3x -1= -2y }\)

OpenStudy (solomonzelman):

Now add the equations, what do you get?

OpenStudy (anonymous):

4x -3 = 0 right ?

OpenStudy (anonymous):

wait 4x -1= 0 ?

OpenStudy (solomonzelman):

yes. 4x-3=0

OpenStudy (solomonzelman):

So, x=?

OpenStudy (anonymous):

3/4 right ?

OpenStudy (anonymous):

so i got the answer (3/4 , 13/8) is that right ?

OpenStudy (solomonzelman):

yes, x=3/4

OpenStudy (solomonzelman):

3x+1=2y 3(3/4)+1=2y 9/4+1=2y 9/4+4/4=2y 13/4=2y y=13/8

OpenStudy (solomonzelman):

very good.

OpenStudy (solomonzelman):

sorry for replying late